OpenStudy (anonymous):

What is 0.4 expressed as a fraction and a percent?

OpenStudy (zehanz):

0.4 = 4/10 = 40/100 = 40 % Fraction 4/10 can also be written as 2/5

OpenStudy (anonymous):

The car Ramon wants to buy is marked down 23%. The original price of the car was $26,000. What is the sale price?